# Emergence of Chimera states in an active fluid medium in the presence of chemical species with logistic growth

### v1

@@ IV Numerical Simulations
Figure 6: Long-lived two-regime transient chimera state. The kymograph displays the spatiotemporal pattern formation from time 100 to 200. The panels on the sides display the time evolution of concentration $A$, at locations in space marked by different symbols in the kymograph. Here the parameters are: $\beta=1.2$, $\alpha=0.1$, $R=1$ and $Pe=13$.
Figure 7: Three-regime transient chimera state. The kymograph displays the spatiotemporal pattern formation from time 0 to 100. The panels on the sides display the time evolution of concentration $A$, at locations in space marked by different symbols in the kymograph. Here the parameters are: $\beta=3$, $\alpha=0.1$, $R=1$ and $Pe=32$.
@@
We solve Eqns. (7)-(9) numerically by using a semi-implicit spectral scheme using periodic boundary conditions. We apply a small localized perturbation to the uniform steady state $(A_{0},I_{0})=(1,1)$, with $v=0$, and follow the system’s response to this perturbation. With no loss of generality with respect to the qualitative features, we consider $L=150$, $\Delta x=L/512$, $\Delta t=0.0001$. Additionally, We have also examined the dynamics with $\Delta x=L/1024,L/2048$ in order to check the robustness of our qualitative results. Specifically, we will demonstrate that a wide variety of spatiotemporal patterns arise from varying strength of the logistic growth term $R$, asymmetry $\beta$ and the Péclet number $Pe$.
@@
The first key observation is that the homogeneous steady state loses stability as the Péclet number $Pe$ increases beyond a critical value, and this critical value is consistent with that obtained through linear stability analysis. When the homogeneous steady state loses stability, a range of patterns emerge. Representative examples of these patterns are displayed in Fig. 4. The patterns are characterized through kymographs, namely the heat map of the concentration of species $A$ in space and time (left panels in Fig. 4). Additionally, the temporal evolution at typical locations in space is characterized by dynamical attractors in the phase space of the concentrations of the two chemical species (right panels in Fig. 4). Typically, for smaller Péclet numbers oscillatory states arise, yielding regular limit cycle attractors in phase space. On the other hand, large Péclet numbers give rise to enhanced irregularity in the spatio-temporal patterns. For instance, in the representative examples in the figure, for $Pe=13$, the emergent space-time patterns are nearly pulsatory, while for $Pe=48$ one obtains a chaotic attractor.
@@
Fig. 5 shows a representative sequence of space-time patterns that emerge for increasing Péclet numbers beyond the regime of regular pulsatory patterns. This illustrative set of patterns reveals the existence of interesting robust and transient chimera states [50] at moderately high values of the Péclet number. On varying initial states, a wide variety of chimera states emerge. These range from long-lived two-regime chimera states to multi-head transient chimera states that emerge spontaneously at uncorrelated locations and times. The specific patterns that emerge are strongly sensitive to initial states, and the persistence of transient chimeras increases with increasing length-scale $L$ of the system. Representative examples of a transient two-regime chimera state and a three-regime chimera state, are shown in Figures 6-7. In the two-regime chimera state (Fig. 6), two types of patterns coexist. We have homogeneous spatial patches where the concentration is nearly stationary for a significant time, with alternating patches having a high and low concentration, i.e. nearly homogeneous stationary patches coexist with a spatial region exhibiting oscillatory space-time patterns. In the three-regime chimera state (Fig. 7), nearly homogeneous steady patches coexist with two separate spatial regions exhibiting distinct oscillatory patterns. The time evolution of concentrations at representative locations in the distinct regimes are explicitly shown in Figs. 6-7. Typically, these transient chimera states appear, disappear and re-appear. This sequence repeats over long times before finally terminating and settling down to the asymptotic oscillatory pattern. In the context of transient chimeras, it is interesting to note that cortical chimeras are often transient and identifying them plays a role in predicting epileptic seizures [51].
@@
In parameter regimes beyond the ones supporting chimera states, almost all initial perturbations give rise to irregular arrhythmic oscillatory spatiotemporal patterns, illustrative examples of which are shown in Fig. 4c-d and Fig. 5d. So when the Péclet number is very small, we obtain stable homogeneous steady states. On increasing $Pe$, chimera states emerge and are found over a considerable range of Péclet numbers. For very high $Pe$, chimera states again disappear, and irregular arrhythmic patterns, akin to spatiotemporal chaos, are commonly observed. So, one can conclude that a moderately high Péclet number is most conducive to chimera-like states.
@@
In the phase diagrams obtained through stability analysis (cf. Figs. 1-2) , the regime of oscillatory instability is most conducive to chimera states. Notice that this regime is most extensive for large $R$, implying that the possibility of observing chimeras is enhanced by larger nonlinearity. This connection with the phase diagram is also consistent with the observation of chimeras at moderate Péclet numbers, but not at high $Pe$ where the oscillatory instability gives way to stationary instability. Lastly the asymmetry parameter $\beta$ needs to be larger than $1$ in order to obtain chimeras. This observation is again linked to the fact that the oscillatory instability exists only for $\beta>1$.
Finally, we observe merging-emerging solitonic structures in the space-time evolution of the concentrations. This class of patterns is illustrated in Fig. 8, which shows the destabilization of the initial near-uniform spatial profile into multi-peak aggregations which evolve through a sequence of emerging and merging events. Such patterns are reminiscent of solitonic defects, where the soliton-like structures emerge, persist for some time and then may disappear, merge with other neighbouring solitonic defects or split into solitonic sub-structures as the system evolves. These patterns are most predominant in the regime of stationary instability. So low asymmetry $\beta$ and high Péclet numbers are most likely to yield them. Such merging and emerging space-time structures are similar to patterns observed in the dynamics of a model for chemotaxis incorporating a logistic cell growth term [35]. Further, such patterns have also been observed in a model system that mimics nonlinear cell-diffusion [52].
@@
Figure 8: (a) Kymograph showing Merging-Emerging space-time dynamics, for parameters $Pe=20$, $R=1$, $\beta=1.2$ and $\alpha=0.1$. The panels (b) and (c) show the spatial variation of concentration $A$ at two instants of time $t=190$ and $t=194$. The emergence and annihilation of concentration peaks are clearly visible in the spatial profile.
V Conclusions
@@
We explored a two component system, where the chemical species interact with the active fluid medium via active stress gradients, and the growth of the fast-diffusing species is modelled by a nonlinear logistic term. We investigated the formation of space-time patterns, both analytically, as well as through extensive numerical simulations. Our key results are summarized below.
@@
First, increasing the Péclet number beyond a critical value leads to the loss of stability of the homogeneous steady state, i.e. instability is induced when the rate of advection is sufficiently greater than the rate of gradient-driven diffusion. Further, the asymmetry of the activator and inhibitor species in our two-component system, as reflected by the parameter $\beta$, also has a crucial effect on the emergent dynamical phases. Increasing asymmetry yields three distinct dynamical phases, while low asymmetry does not yield the oscillatory instability. Lastly, most interestingly, we demonstrate that nonlinear growth aids the stability of the homogeneous steady state. This counter-intuitive phenomenon is evident from the enlarged area of stability of the homogeneous steady state in the phase diagrams, for larger scaled nonlinear growth parameter $R$. Specifically we also show that the critical Péclet number for the onset of instability rises approximately linearly with $R$, leading to an expansion of the homogeneous steady state region in parameter space. Our analytical results explicitly showing the emergence of different dynamical phases are completely corroborated by numerical simulations.
@@
The second set of significant results arise in the region of parameter space where the homogeneous steady state loses stability. This region is explored extensively numerically, as well as through the determination of the existence of complex eigenvalues for certain modes in the perturbed system indicating oscillatory instabilities. We find the emergence of diverse classes of patterns in this regime. These range from regular oscillatory space-time patterns, to merging-emerging multi-peak aggregations reminiscent of solitonic defect-like structures, and irregular space-time evolution characterized by chaotic attractors. Most interestingly, we find the emergence of chimera states, both long-lived transient, as well as robust, and these chimeras are predominant in the region of oscillatory instability. Now, the dispersion curves for different modes of the perturbed system obtained through our linear stability analysis show that the region of oscillatory instability is enlarged in the presence of nonlinear growth. This implies the following important result: nonlinear growth enhances the probability of observing chimera states. So these results can potentially aid experimentalists, as they can focus their search for chimera states on classes of systems that are more prone to such patterns.

### v2

@@ 4 Numerical Simulations
We solve Eqns. (7)-(9) numerically by using a semi-implicit spectral scheme using periodic boundary conditions. We apply a small localized perturbation to the uniform steady state $(A_{0},I_{0})=(1,1)$, with $v=0$, and follow the system’s response to this perturbation. Such localized perturbations help us track the spread of a small disturbance in the system and provide a method to gauge the response of the spatially extended system. With no loss of generality with respect to the qualitative features, we consider $L=150$, $\Delta x=L/512$, $\Delta t=0.0001$. Additionally, we have also examined the dynamics with $\Delta x=L/1024,L/2048$ in order to check the robustness of our qualitative results. We find that the emergent patterns are qualitatively robust with respect to system size. Further note that spatiotemporal patterns qualitatively similar to those displayed in our figures emerge from a large sample of random initial conditions at the specified parameter values.
@@
Now we will demonstrate the wide variety of spatiotemporal patterns that arise from varying strength of the logistic growth term $R$, asymmetry $\beta$ and the Péclet number $Pe$. The first key observation is that the homogeneous steady state loses stability as the Péclet number $Pe$ increases beyond a critical value, and this critical value is consistent with that obtained through linear stability analysis. When the homogeneous steady state loses stability, a range of patterns emerge. Representative examples of these patterns are displayed in Fig. 4. The patterns are characterized through kymographs, namely the heat map of the concentration of species $A$ in space and time (left panels in Fig. 4). Additionally, the temporal evolution at typical locations in space is characterized by dynamical attractors in the phase space of the concentrations of the two chemical species (right panels in Fig. 4). Typically, for smaller Péclet numbers oscillatory states arise, yielding regular limit cycle attractors in phase space. On the other hand, large Péclet numbers give rise to enhanced irregularity in the spatio-temporal patterns. For instance, in the representative examples in the figure, for $Pe=13$, the emergent space-time patterns are nearly pulsatory. On the other hand, for higher Péclet numbers such as $Pe=48$, one obtains a state characterized by long-term aperiodic temporal dynamics. This dynamics gives rise to a chaotic attractor in phase space, exhibiting a complex irregular geometric structure, in stark contrast to the regular simple limit cycle for lower Péclet numbers.
Figure 4: [(a),(e)] Kymographs displaying the spatiotemporal evolution of concentration AA; [(b),(f)] dynamical attractors in the phase space of AIA-I concentrations at a typical location in space; [(c),(g)] snapshot of the spatial profile; and [(d),(h)] power spectrum (PSD) of the time series of $A$ at a site with spatial PSDs shown in the insets. Upper row corresponds to the parameters $Pe=13$, $\beta=1.5$, $R=1$, $\alpha=0.1$, and lower row corresponds to the $Pe=48$, $\beta=3$, $R=1$, $\alpha=0.1$. (b) and (c) display regular limit cycle behaviour and a regular oscillatory spatial profile. (f) and (g) display a chaotic attractor corresponding to the irregular spatial profile. Note that the system evolves to such attractors from a wide range of initial states, after transience.
Figure 5: Emergence of an illustrative sequence of space-time patterns, ranging from transient Chimera states to irregular arrhythmic spatiotemporal patterns, in a system with $R=1$, $\alpha=0.1$, $\beta=3$, for increasing Péclet numbers: (a) $Pe=27$, (b) $Pe=35$, (c) Pe=39Pe=39,(d) $Pe=48$.
Figure 6: Long-lived two-regime transient chimera state. (a) Kymograph displaying the spatio-temporal pattern. (b) and (c) show the time evolution of concentration AA, at locations in space marked by different symbols in the kymograph. (d) displays the mean value $\mu$ of the concentration of $A$ at different sites across the lattice, showing markedly different values in the distinct spatial domains, thereby serving as a good quantitative measure of the chimera state. The parameters here are: $\beta=1.2$, $\alpha=0.1$, $R=1$ and $Pe=13$.
Figure 7: Three-regime transient chimera state. The kymograph displays the spatiotemporal pattern formation from time 0 to 100. The panels on the sides display the time evolution of concentration AA at locations in space marked by different symbols in the kymograph. Here the parameters are: $\beta=3$, $\alpha=0.1$, $R=1$ and $Pe=32$. Note that the asymptotic state of the system is pulsatory, displaying regular oscillations in time and space.
@@
Figure 5 shows a representative sequence of space-time patterns that emerge for increasing Péclet numbers beyond the regime of regular pulsatory patterns. This illustrative set of patterns reveals the existence of interesting transient chimera states for moderately high Péclet numbers. Typically, the transient chimera states are non-stationary, and they appear, disappear and re-appear. This sequence repeats over long times before finally terminating and settling down to the asymptotic oscillatory pattern. In the context of transient chimeras, it is interesting to note that chimeras in the human cortex are often transient, and identifying them plays a role in predicting epileptic seizures [57].
Now, a wide variety of chimera states emerge dependent on initial states, for large enough systems, with the persistence of transient chimeras increasing with increasing length-scale $L$ of the system. A representative example of a transient two-regime chimera state is shown in Fig. 6. Here, two types of patterns coexist. We have homogeneous spatial patches where the concentration is nearly stationary for significant times, co-existing with alternating patches having a high and low concentration, i.e. nearly homogeneous stationary patches coexist with a spatial region exhibiting oscillatory space-time patterns. An example of a three-regime chimera state is displayed in Fig. 7. Here, nearly homogeneous steady patches coexist with two separate spatial regions exhibiting distinct oscillatory patterns. The time evolution of concentrations at representative locations in the distinct regimes is explicitly shown in Figs. 6-7. Notice that in the multi-domain chimera displayed in Fig. 7, the dynamics in the domains (b) and (e), while qualitatively similar, exhibit different degrees of regularity and frequencies of the oscillatory behaviour which can be discerned from the dominant frequencies in the Fourier spectra and the level of the noisy background vis-a-vis the peaks. Further, notice that the oscillatory spatial domains are separated by domains of temporal stationarity (evident as dark vertical domains in the kymographs), and these domains are also of significantly different spatial extents. Interestingly, in the coherent regimes of the chimaera states, the net flux and the magnitude of the growth term are close to zero, and so these regimes behave like a spatial barrier, with soliton-like waves moving, colliding and passing through each other, but reflected back at the boundary of the coherent regimes. So these waves are trapped within a domain between the two coherent regimes, forming a distinct spatial domain with temporal behaviour very different from the coherent domain, i.e. the chimera states are characterized by the coexistence of irregular motion of soliton-like waves and regions of coherence with almost zero concentration. Additionally, we find the emergence of robust chimera states, as shown in Fig. 8.
Figure 8: Kymographs displaying robust multi-head chimera states, for parameter values $Pe=21$, $\beta=2$, $R=1$, $\alpha=0.1$. Here time ranges from $400$ to $1100$, across the four panels, from left to right.
Figure 9: Dependence of the strength of incoherence $S$ (see text for definition) on the Péclet number. Here $\beta=2$, $\alpha=0.1$, and $R=1$ (blue circles) and $R=0.25$ (red diamonds). The quantity $S$ takes values ranging from 0 (characteristic of coherent states) to $1$ (characteristic of completely incoherent states), and takes values within the range $(0,1)$ for chimera states. The emergence of chimera states after $Pe\sim 15$ is clearly evident for $R=1$, while it does not occur for $R=0.25$.
@@
In parameter regimes beyond those supporting chimera states, almost all initial perturbations give rise to irregular arrhythmic oscillatory spatio-temporal patterns, illustrative examples shown in Figs. 4 and Fig. 5d. So when the Péclet number is very small, we obtain stable homogeneous steady states. As $Pe$ increases, chimera states emerge and are found over a considerable range of Péclet numbers. For very high $Pe$, chimera states again disappear, and irregular arrhythmic patterns, akin to spatiotemporal chaos, are commonly observed. So, one can conclude that a moderately high Péclet number is most conducive to chimera-like states.
@@
In the phase diagrams obtained through stability analysis (cf. Figs. 1-2), the regime of oscillatory instability is most conducive to chimera states. Notice that this regime is most extensive for large $R$, implying that the possibility of observing chimeras is enhanced by larger nonlinearity. This connection with the phase diagram is also consistent with the observation of chimeras at moderate Péclet numbers but not at high $Pe$, where the oscillatory instability gives way to stationary instability. Lastly, the asymmetry parameter $\beta$ needs to be larger than $1$ to obtain chimeras. This observation is again linked to the fact that the oscillatory instability exists only for $\beta>1$.
@@
Figure 10: Merging-emerging spatiotemporal dynamics. (a) Kymograph showing merging-emerging dynamics, for parameters $Pe=20$, $R=1$, $\beta=1.2$ and $\alpha=0.1$. Panels (b) and (c) show the spatial variation of concentration AA at two instants of time $t=190$ and $t=194$. The emergence and annihilation of concentration peaks are clearly visible in the spatial profile.
To further quantify the prevalence of chimeras, we have computed the strength of incoherence which serves as a good order parameter to capture the existence of chimeras and provides a measure to characterize chimera states in complex systems. This quantity is defined as [58]: $S=1-\frac{\sum_{x}H(\bar{A})}{N},H(\bar{A})=\Theta[\delta-\bar{A}]$, where $\bar{A}=\langle A(x)\rangle-\langle A_{c}\rangle$. Here, $\bar{A}$ measures the deviation of species densities at $x$ from the coherent subpopulation, $\langle A_{c}\rangle$ represents the average amplitude of the coherent subpopulation, and $\delta$ is a threshold value. The quantity $S$ takes values ranging from 0 (characteristic of coherent states) to $1$ (characteristic of completely incoherent states), and takes values within the range $(0,1)$ for chimera states, indicating the coexistence or coherence and incoherence. So $S$ can serve as an “order parameter” indicating the occurrence of chimeras. Fig. 9 shows this quantity over a range of Péclet numbers. Clearly, $S=0$ when $R$ is low (specifically $R=0.25$ in the figure). On the other hand, chimeras occur in a window of Péclet number when $R$ is sufficiently large (specifically $R=1$), where we have a homogeneous steady state till $Pe\sim 11$ (with $S=0$), a regular pulsatory state for $Pe\sim 12-15$ (with $S=0$), and chimeras beyond $Pe\sim 15$ (with $0<S<1$).
Finally, we observe merging-emerging solitonic structures in the space-time evolution of the concentrations. Such patterns involve the spontaneous emergence of localized concentration peaks (referred to as “emerging”) and the coalescing of two existing localized concentration peaks (referred to as “merging”). This class of patterns is illustrated in Fig. 10, which shows the destabilization of the initial near-uniform spatial profile into multi-peak aggregations, which evolve through a sequence of emerging and merging events. Such patterns are reminiscent of solitonic defects, where the soliton-like structures emerge, persist for some time and then may disappear, merge with other neighbouring solitonic defects or split into solitonic sub-structures as the system evolves. These patterns are most predominant in the regime of stationary instability. So low asymmetry $\beta$ and high Péclet numbers are most likely to yield them. Such merging and emerging space-time structures are similar to patterns observed in the dynamics of a model for chemotaxis incorporating a logistic cell growth term [43]. Further, such patterns have also been observed in a model system that mimics nonlinear cell-diffusion [59].
In this study, we considered the widely used periodic boundary conditions. Prototypically, periodic boundary conditions are chosen to approximate large systems and eliminate edge effects in simulations, and for these reasons, they have been broadly employed in the study of model systems mimicking dynamics within a fluid environment or membrane-like structure. Of course, the influence of different boundary conditions on emergent patterns is an open problem of considerable relevance and holds much potential for future investigations.
5 Two-dimensional Systems
We have extended our simulations to two dimensions. The governing equations are now expressed as
$\partial_{t}A=-\nabla\cdot(\mathbf{\mathbf{v}}A)+D\nabla^{2}A+rA(1-\frac{A}{K})\\$ (14)
$\partial_{t}I=-\nabla\cdot(\mathbf{\mathbf{v}}I)+\alpha D\nabla^{2}I\\$ (15)
$\nabla\cdot\mathbf{\sigma}=\gamma\mathbf{v}$ (16)
$\sigma=\sigma_{p}+\sigma_{a}\mathbb{I}$ (17)
$\sigma_{p}=[\nabla\mathbf{v}+(\nabla\mathbf{v})^{T}-\frac{2}{d}(\nabla\cdot\mathbf{v})\mathbb{I}]+\eta_{v}(\nabla\cdot\mathbf{v})\mathbb{I}$ (18)
$\sigma_{a}=\sigma_{0}f(A,I)$ (19)
Here, $\eta$ and $\eta_{v}$ represent the shear and bulk viscosity, $d$ is the dimension (here, $d=2$). Specifically, we choose $\eta_{v}$ = $3\eta$ [60, 61] and use a semi-implicit spectral scheme to simulate the dynamics given by the equations above. Pattern formation is explored mostly in the square domain, with periodic boundary conditions, under a small local perturbation. In the dimensionless form, the equations are given below:
$\partial_{t}A=-\nabla\cdot(\mathbf{\mathbf{v}}A)+\nabla^{2}A+RA(1-A)\\$ (20)
$\partial_{t}I=-\nabla\cdot(\mathbf{\mathbf{v}}I)+\alpha\nabla^{2}I\\$ (21)
$\kappa\partial^{2}_{x}u+\epsilon\frac{\partial^{2}v}{\partial x\partial y}+Pe\frac{\partial f}{\partial x}+\partial^{2}_{y}u=u\\$ (22)
$\kappa\partial^{2}_{y}v+\epsilon\frac{\partial^{2}u}{\partial x\partial y}+Pe\frac{\partial f}{\partial y}+\partial^{2}_{x}v=v\\$ (23)
where, $\kappa=1+\epsilon$ with $\epsilon=\frac{\eta_{v}}{\eta}$. Also, $\mathbf{v}=(u,v)$.
As in the $1$-dimensional case detailed earlier, we observe homogeneous steady states and regular pulsatory patterns, as well as stationary inhomogeneous patterns and symmetry-breaking amplitude chimera states (Figs. 11-13). In particular, we present a few illustrative spatio-temporal patterns in Fig. 11, showing the emergence of diverse pulsatory patterns. For instance, for a square domain with a perturbation at a random site, we obtain periodic circular patterns with a circle that alternate between high and low concentrations, evident as periodically changing bright and dark circles in the heat map of the domain. If perturbations are applied at two random sites, alternating bright-dark circular structures emerge. We have also explored pattern formation in a rectangular domain. We find interesting pulsatory band-like structures (see Fig. 12), reminiscent of splay patterns [62], as well as chimera states with two regimes with distinct amplitudes (see Fig. 13). Note that we have explored system sizes an order of magnitude larger than those displayed in these figures, and we find that qualitatively similar patterns emerge for larger systems as well.
So it is apparent that two-dimensional active fluid systems offer a rich repertoire of spatiotemporal patterns, ranging from homogeneous steady states to chimera states. Our results here can then trigger future avenues of exploration of chimeras in active fluid systems in domains of different geometries under different boundary conditions and perturbations.
Figure 11: Snapshots of the $2$-dimensional domain, at different instants of time: (top row) Pulsatory spatiotemporal patterns under perturbation at one central site. (Bottom row) Two distinct pulsatory spatial domains, with alternating high and low concentrations, under perturbation at two random sites. Here the heat map show the concentration of $A$, and parameter values are $Pe=40$, $\beta=3$ , $R=1$ and $\alpha=0.1$. The step-size $\Delta t$ = $0.001$ and $\Delta x$ = $0.03$ in the simulation, with $L_{x}=L_{y}=5$.
Figure 12: Snapshots of a 22-dimensional rectangular domain, with $L_{x}=20,L_{y}=5$, at different instants of time, showing pulsatory band-like spatiotemporal patterns. Here the system displays synchronized vertical domains, and splits into multi-clusters along the horizontal axis.
Figure 13: Chimera state in two-dimensional systems: the kymograph shows a horizontal section in the square lattice for $Pe=80$, $\beta=3$, $R=1$, $\alpha=0.1$, $L_{x}=L_{y}=5$ (Left) and $Pe=50$, $\beta=3$, $R=1$, $\alpha=0.1$, $L_{x}=L_{y}=15$ (Right), alongside time series at two representative spatial locations placed in the two distinct domains, as indicated by different symbols on the kymograph.
6 Conclusions
@@
We explored a two-component system, where the chemical species interact with the active fluid medium via active stress gradients, and the growth of the fast-diffusing species is modelled by a nonlinear logistic term. We investigated the formation of space-time patterns, both analytically and through extensive numerical simulations. Our key results are summarized below.
@@
First, increasing the Péclet number beyond a critical value leads to the loss of stability of the homogeneous steady state, i.e. instability is induced when the rate of advection is sufficiently greater than the rate of gradient-driven diffusion. Further, the asymmetry of the activator and inhibitor species in our two-component system, as reflected by the parameter $\beta$, also has a crucial effect on the emergent dynamical phases. Increasing asymmetry yields three distinct dynamical phases, while low asymmetry does not yield oscillatory instability. Lastly, most interestingly, we demonstrate that nonlinear growth aids the stability of the homogeneous steady state. This counter-intuitive phenomenon is evident from the enlarged area of stability of the homogeneous steady state in the phase diagrams for larger-scaled nonlinear growth parameter $R$. Specifically, we also show that the critical Péclet number for the onset of instability rises approximately linearly with $R$, leading to an expansion of the homogeneous steady state region in parameter space. Our analytical results explicitly show the emergence of different dynamic phases, which are entirely corroborated by numerical simulations.
@@
The second set of significant results arises in the region of parameter space where the homogeneous steady state loses stability. This region is explored extensively numerically, and complex eigenvalues for certain modes in the perturbed system are determined, indicating oscillatory instabilities. We find the emergence of diverse classes of patterns in this regime. These range from regular oscillatory space-time patterns to merging-emerging multi-peak aggregations reminiscent of solitonic defect-like structures and irregular space-time evolution characterized by chaotic attractors. Most interestingly, we find the emergence of chimera states, both long-lived transient and robust, and these chimeras are predominant in the region of oscillatory instability. Now, the dispersion curves for different modes of the perturbed system obtained through our linear stability analysis show that the region of oscillatory instability is enlarged in the presence of nonlinear growth. This implies the following important result: nonlinear growth enhances the probability of observing chimera states. These results can potentially aid experimentalists, as they can focus their search for chimera states on classes of systems more prone to such patterns.
Appendix: Reduced-Order Analysis - Galerkin Projection and SVD
To complement the direct numerical simulations and linear stability analysis, we perform a reduced-order study using two approaches: (i) Galerkin mode truncation and (ii) Singular Value Decomposition (SVD). While the Galerkin method provides a low-dimensional approximation of the governing PDEs by projecting them onto a truncated set of orthogonal modes, SVD offers a data-driven decomposition of the spatio-temporal patterns, revealing the hierarchy of dominant modes.
Figure 14: 22-mode coupling results: upper row displays the time series of $A_{2}$, and the lower row displays the phase portraits of the temporal coefficients in the space of $A_{2}-I_{2}$, for the parameters are (left to right): $Pe=20,40,50,58,68$, with $\beta=1.75$, $R=1$, $\alpha=0.1$, $L=50$.
Figure 15: (a) Kymograph of Pulsatory patterns of the perturbation ($A_{p}$)from $Pe=27,\beta=3,R=1,\alpha=0.1,L=50$ and (b) image represents the amplitude Chimera states of the perturbation($A_{p}$) for AA with $Pe=33,\beta=3,R=1,\alpha=0.1,L=50$ (obtained from Mode 2 coupling). (c) Kymograph of AA for $Pe=11,\beta=2,R=1,\alpha=0.1,L=25$ (Chimera states) and (d) multi-valued Stationary states kymograph for AA for $Pe=12,\beta=2,R=1,L=40$ (obtained from 33 mode truncation analysis).
A. Galerkin Mode Truncation Analysis
We first apply Galerkin projection [63] to approximate the temporal evolution of the dominant modes. The concentration fields are expanded around the homogeneous steady state:
$A=A_{0}+A_{p}(x,t),\quad I=I_{0}+I_{p}(x,t),\quad v=v_{0}+v_{p}(x,t).$ (24)
Cosine functions are chosen as basis functions due to their orthogonality under periodic boundary conditions. Considering the first two modes, the perturbations are written as:
$\epsilon(x,t)=\epsilon_{1}(t)\cos\left(\frac{2\pi x}{L}\right)+\epsilon_{2}(t)\cos\left(\frac{4\pi x}{L}\right),$ (25)
$A_{p}(x,t)=A_{1}(t)\cos\left(\frac{2\pi x}{L}\right)+A_{2}(t)\cos\left(\frac{4\pi x}{L}\right),$ (26)
$I_{p}(x,t)=I_{1}(t)\cos\left(\frac{2\pi x}{L}\right)+I_{2}(t)\cos\left(\frac{4\pi x}{L}\right),$ (27)
where $\epsilon=\partial v/\partial x$ denotes the strain rate. For simplicity, we use a linearized active stress function:
$f(A,I)=(1+\beta)\frac{A}{A_{S}}+(1-\beta)\frac{I}{I_{S}}.$ (28)
Using orthogonality, the evolution equations for the amplitudes $A_{i},I_{i}$ are derived as:
$\epsilon_{1}=\frac{-Pe\,k^{2}\left[(1+\beta)\frac{A_{1}}{A_{S}}+(1-\beta)\frac{I_{1}}{I_{S}}\right]}{1+k^{2}},$ (29)
$\epsilon_{2}=\frac{-4Pe\,k^{2}\left[(1+\beta)\frac{A_{2}}{A_{S}}+(1-\beta)\frac{I_{2}}{I_{S}}\right]}{1+4k^{2}},$ (30)
$\frac{dA_{1}}{dt}=-\epsilon_{1}+\frac{3A_{1}\epsilon_{2}}{4}+\frac{3A_{2}\epsilon_{1}}{2}-k^{2}A_{1}-RA_{1}-RA_{1}A_{2},$ (31)
$\frac{dA_{2}}{dt}=-\epsilon_{2}-A_{1}\epsilon_{1}-k^{2}A_{2}-RA_{2}-\frac{RA_{1}^{2}}{2},$ (32)
$\frac{dI_{1}}{dt}=-\epsilon_{1}+\frac{3I_{1}\epsilon_{2}}{4}+\frac{3I_{2}\epsilon_{1}}{2}-\alpha k^{2}I_{1},$ (33)
$\frac{dI_{2}}{dt}=-\epsilon_{2}-I_{1}\epsilon_{1}-\alpha k^{2}I_{2}.$ (34)
This two-mode truncation captures key qualitative features of the system, including steady states, oscillations, and even more interestingly, amplitude chimera states (see Figs. 14-15). However, it is limited in scope and fails to preserve certain intrinsic properties of the full PDEs, such as global conservation and boundedness [64]. For instance, unphysical negative concentration values may arise in the reduced model. Extending the analysis to three modes (yielding six coupled ODEs) mitigates such artifacts and provides better agreement with the full numerical solutions. For example, chimera states are observed for $Pe=11,\beta=2,R=1,\alpha=0.1,L=25$ in the three-mode model (see Fig. 15), consistent with direct PDE simulations where chimera states emerge at $Pe=17$. Phase portraits of the amplitude variables $(A_{1},A_{2},A_{3},I_{1},I_{2},I_{3})$ exhibit a mix of regular (limit-cycle-like) and complex (folding and stretching) structures, indicating quasi-periodic or aperiodic behavior and hinting at the presence of chimera-like states (see bottom row of Fig. 14).
Figure 16: (a-e) Reconstructed kymograph with first 33 dominant modes obtained from SVD, three optimal basis functions and relative strength of each basis modes for $Pe=13,\beta=1.2,R=1,\alpha=0.1$. (f-j) Reconstructed kymograph with first 33 dominant modes obtained from SVD, three optimal basis functions and relative strength of each basis modes for $Pe=15,\beta=3,R=1,\alpha=0.1$.
Figure 17: 3D attractors of $A_{1},A_{2},A_{3}$ and $I_{1},I_{2},I_{3}$ corresponding to 3 mode coupling, for $Pe=11,\beta=2,R=1,\alpha=0.1,L=25$.
B. Singular Value Decomposition (SVD)
While the Galerkin approach imposes a predefined basis, the SVD provides a data-driven decomposition of the spatio-temporal fields. We apply SVD [65] to the simulation data to identify the optimal orthogonal modes and their relative strengths. For a data matrix $A(x,t)$, the SVD is expressed as:
$A(x,t)=\sum_{j=1}^{N}\sigma_{j}\,\phi_{j}(x)\,\psi_{j}(t),$ (35)
where $\phi_{j}(x)$ and $\psi_{j}(t)$ are spatial and temporal modes, respectively, and $\sigma_{j}$ are singular values. The normalized energy contribution of mode $j$ is:
$Y_{j}=\frac{\sigma_{j}^{2}}{\sum_{i=1}^{N}\sigma_{i}^{2}}.$ (36)
The SVD analysis reveals that in regimes of simple pulsatory dynamics (e.g., $Pe=15,\beta=3,R=1,\alpha=0.1$), the first mode dominates ($Y_{1}\approx 1$) while higher modes are negligible. In contrast, for chimera states (e.g., $Pe=13,\beta=1.2,R=1,\alpha=0.1$), the first three modes capture over $97\%$ of the energy, with $\phi_{1}$ representing coherent spatial regions and $\phi_{2},\phi_{3}$ capturing oscillatory variations. For more complex regimes such as spatio-temporal chaos or merging-emerging patterns (e.g., $Pe=18,\beta=1.2,R=1,\alpha=0.1$), up to $\sim 200$ modes may be required for accurate reconstruction. Here, the energy spectrum $Y_{j}$ decays slowly, indicating strong contributions from higher-order modes (see Fig. 16).
Examining the temporal coefficients $(A_{1},A_{2},A_{3})$ and their phase portraits provides additional insights: pulsatory states display regular closed loops (limit cycles), while chimera and chaotic states show folding and stretching, reflecting the coexistence of coherent and incoherent dynamics (see Fig. 17). Thus, SVD complements the Galerkin analysis by providing a compact, data-driven framework to dissect the hierarchical structure of spatio-temporal patterns.